OpenStudy (mathteacher1729):

Here is a challenging counting / combinatorics problem: How many 5-letter permutations are there of the word "MATHEMATICS" ? Explain your solution. :)

OpenStudy (anonymous):

guess the answer is 11 ! /8 assuming you cannot replace letters

OpenStudy (anonymous):

well that is silly, not 11! /8 but rather \[\frac{11\times 10\times 9\times 8\times 7}{8}\]

OpenStudy (anonymous):

interpreting the question to mean essentially you have 11 tiles with those letters on them like in scrabble.

OpenStudy (anonymous):

wouldn't it be 11*10*9*8*7?

OpenStudy (anonymous):

well it would be but you cannot tell the two m's apart in your counting mathm counts twice

OpenStudy (anonymous):

likewise for A and T which is you divide by 8

OpenStudy (anonymous):

what about the A and T's =- oh i see

OpenStudy (anonymous):

2 * 2 * 2

OpenStudy (anonymous):

yeah that explains the 8 in the denominator. if you want to write something really stupid you could put \[\frac{11!}{(11-5)!2!2!2!}\] and make it look fancy

OpenStudy (anonymous):

right

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i beleive you it is early

OpenStudy (mathteacher1729):

@Dangol -- no, sorry it is is not that. Your answer is also too large.

OpenStudy (anonymous):

i got 6930 is that wrong?

OpenStudy (mathteacher1729):

HINT: There are three cases we must consider, with respect to repetitions. 1) No repeated letters 2) 1 pair of repeated letters 3) 2 pairs of repeated letters

OpenStudy (mathteacher1729):

Sorry, Satellite, your answers are too SMALL., my mistake. :(

OpenStudy (mathteacher1729):

The answer is between 55,440 and 6,930. I double-checked.

OpenStudy (anonymous):

give me a minute see if i can find my mistake. "minute" means ...

OpenStudy (anonymous):

ok i still don't see it so let me tell you what i was thinking and you can tell me why i am wrong. oh wait i think i see the problem.

OpenStudy (anonymous):

i shouldn't divide the whole thing by 8. only divide the ones that have 2 m's, 2 a's or to t's by 2 yikes this is going to take me forever

OpenStudy (anonymous):

oh and then of course there is the possibility that it has 2 a's and 2 t's etc

OpenStudy (anonymous):

give me something to think about while i go run my saturday errands. don't tell me the answer, but is there a snap way of doing this of doing this?

OpenStudy (mathteacher1729):

You have to take the three cases individually and add them: 1) How many 5-letter permutations of MATHEMATICS are there, where there are no repetitions. 2) How many 5-letter permutations of MATHEMATICS are there, where there are one pair of repeated letters? 3) How many 5-letter permutations of MATHEMATICS are there, where there are two pairs of repeated letters? Part 1 is the easiest to answer.

OpenStudy (anonymous):

got it! i mean i have the idea, not the answer. thnx

OpenStudy (zarkon):

a quick calculation gives me 13560